Find the general solution of differential equation (d^2-4d+2)y=0


1
Expert's answer
2021-05-07T13:58:02-0400

We have given the differential equation:


"(D^2-4D+2)y = 0"


Its auxiliary equation can be written as :


"m^2-4m+2 = 0"


"m = \dfrac{4 \pm 4i}{2}"


"m = 2 \pm 2i"


Hence, its solution can be written as,


"y = e^{2x}(C_1cos2x+C_2sin2x)"